# Label Smoothing for Enhanced Text Sentiment Classification

### 11 Dec 2023 v1

@@ Abstract
Label smoothing is a widely used technique in various domains, such as image classification and speech recognition, known for effectively combating model overfitting. However, there is few research on its application to text sentiment classification. To fill in the gap, this study investigates the implementation of label smoothing for sentiment classification by utilizing different levels of smoothing. The primary objective is to enhance sentiment classification accuracy by transforming discrete labels into smoothed label distributions. Through extensive experiments, we demonstrate the superior performance of label smoothing in text sentiment classification tasks across eight diverse datasets and deep learning architectures: TextCNN, BERT, and RoBERTa, under two learning schemes: training from scratch and fine-tuning.
**footnotetext: The corresponding author is Shijing Si, shijing.si@outlook.com.
1 Introduction
@@
Text sentiment classification is to identify and extract emotional tendencies (such as positive, negative or neutral) from text, which provides important application scenarios and challenges for the development of natural language processing (NLP)(Li et al.,, 2022). Many methods have been proposed for this task, for instance, convolutional neural networks (CNN) (Zhang et al.,, 2015), and recurrent neural networks (RNN) with attention mechanism (Zhang et al.,, 2016; Chen et al.,, 2017). Transformer-based models (Jiang et al.,, 2021; Chang et al.,, 2020) have excelled in sentiment classification tasks, achieving state-of-the-art performance (Liu et al., 2022a, ; Liu et al., 2022c, ).
@@
Label smoothing (LS) proceeds by using soft targets that are a weighted average of the hard targets and the uniform distribution over labels (Müller et al.,, 2019; Lienen and Hüllermeier,, 2021). This technique lessens the disparity between the top probability estimate and the remaining ones, thereby acting as a barrier to the model from generating extremely confident predictions, consequently decreasing the model’s likelihood of becoming excessively tailored to the training data (Lukasik et al.,, 2020; Gao et al.,, 2020). This also fosters generalization, leading to better performance on unseen data (Chen et al.,, 2020; Cui et al.,, 2022). Furthermore, LS can also mitigate the impact of noisy labels on the training process(Li et al.,, 2020).
@@
LS has achieved widespread success in NLP (Chandrasegaran et al.,, 2022; Liu et al., 2022b, ) by introducing soft targets, which allows the model to optimize towards a more flexible direction during training(Haque et al.,, 2023; Pan et al.,, 2022). However, the research on LS for text sentiment classification is still limited. For the task of sentiment classification, emotion is not an absolute and discrete concept (Pang et al.,, 2002; Onan,, 2022), but there is certain fuzziness and continuity (Hasib et al.,, 2021). LS can take account of this ambiguity, allowing the model to learn more balanced relative probabilities between different emotions.
@@
In order to validate whether LS can benefit sentiment classification, we conduct extensive experiments on three widely used deep neural network architectures: TextCNN, BERT, and RoBERTa, under two learning schemes: training from scratch and fine-tuning. We investigate the performance of LS on six three-class datasets and two binary-class datasets. A comparison with the baseline models is also conducted. By varying the smoothing parameter, LS approach can outperform the baseline method across all architectures.
@@
The remaining sections of this paper are organized as follows. We first review the related work in Section 2. Then, in Section 3, we propose the application of LS in text sentiment classification, including the definition of the text LS learning problem, the theoretical foundations of text LS learning and the selected deep learning architecture. Subsequently, in Section 4, we present the experimental results. Finally, Section 5 provides a discussion, and Section 6 concludes the paper.
2 Related Work
@@
Text sentiment classification is one special type of text classification Si et al., (2020); Kowsari et al., (2019), as the labels are ordinal Si et al., (2022). Since the 2010s, deep learning models (Zulqarnain et al.,, 2020; Liu et al.,, 2016), which automatically extract meaningful representations for text without the need for manual rule and feature design(Minaee et al.,, 2021), are increasingly utilized for text classification.
@@
LS is firstly introduced in the field of computer vision (CV) and has achieved success in various visual recognition tasks(Gao et al.,, 2017; Müller et al.,, 2019). Later, the method is shown to be effective in Machine Translation(MT)(Gao et al.,, 2020). Moreover, it also has applications in sentiment classification(Luo et al.,, 2022, 2021) and Named Entity Recognition(NER)(Zhu and Li,, 2022; Yu et al.,, 2022), improving model calibration and bringing flatter neural minima.
@@
Meanwhile, researchers have started to apply LS to text sentiment classification. One attempt is to add LS to the loss function to enhance the performance of the model (Wang et al., 2022a, ). Similarly, another group add LS to loss function and perform emotion classification on the adaptive fusion features obtained (Yan et al.,, 2023). Also, properties of LS and its adversarial variants are studied, showing LS can enhance the adversarial robustness of the model(Yang et al.,, 2022). Yan et al., (2023) proposes a new cyclic smoothing labeling technique for handling the periodicity of angles and increasing error tolerance for adjacent angles. They also designed a densely-coded tag that greatly reduced the length of the code. Wu et al., (2022) propose an efficient data augmentation method, termed text smoothing, by converting a sentence from its one-hot representation to a controllable smoothed representation and show that text smoothing outperforms various mainstream data augmentation methods by a substantial margin.
@@
Although LS is a common label processing technique, there are still some problems and challenges that need further study. Examples include comparison of effects across different data sets, tasks, and model structures, comparison with other regularization methods, etc(Huang et al.,, 2020). Studying LS can explore these problems and promote the further development of the field of LS. In this paper, we will focus on how to exploit LS in text classification, and a discussion about its superiority and principle will also be given.
3 LS Method on Texts Classification
3.1 Label Smooth Processing
@@
Given a tokenized input $(x_{1},x_{2},x_{3},\ldots,x_{n})$ and a set of labels $(y_{1},y_{2},y_{3},\ldots,y_{k})$, where $n$ is the length of input and $k$ is the number of categories for classification. Thus, $D_{i}=(d_{x_{i}}^{y_{1}},d_{x_{i}}^{y_{2}},d_{x_{i}}^{y_{3}},\ldots,d_{x_{i}}^{y_{k}})$ describes the extent to which the instance $x_{i}$ belongs to a label $y_{k}$. We expect to obtain a final classification result $d_{i}$ which is typically the maximum value of the label distribution $D_{i}$ obtained by applying a softmax normalization to the last layer of the deep neural network. The label corresponding to the maximum value represents the category to which this text segment belongs.
Figure 1: Example of single-label, multi-label and LS
Take the instance in Figure 1 as an example, in single-label conditions, $d_{x}^{y_{2}}=1$ and the rest labels all equal to 00. In multi-label conditions, $d_{x}^{y_{2}}=d_{x}^{y_{5}}=0.5$ and the rest labels are 00. Both two methods only represents part of the labels and ignore the others. Finally, in LS, all of the labels are given concentrations and they also add up to $1$. Meanwhile, the dominance of The most significant label $d_{x}^{y_{2}}$ should be maintained to ensure that the model learns the primary features for sentiment classification.
@@
Due to the continuity of label distribution, it has a larger labeling space, a broader range of expression, and therefore can provide greater flexibility in the learning process. Specifically, in order to convert hard labels to soft label distributions, we use LS method, performing a linear transformation on the original label distribution $D_{i}=(d_{x_{i}}^{y_{1}},d_{x_{i}}^{y_{2}},d_{x_{i}}^{y_{3}},\ldots,d_{x_{i}}^{y_{k}})$ to obtain $D_{i}^{{}^{\prime}}=(1-k\lambda)D+\lambda,\lambda\in[0,1]$.
@@
Suppose that $\theta$ is the parameters of the models waiting to be trained, and $P(y_{k}|x_{i};\theta)=p(y_{1}|x_{i};\theta),p(y_{2}|x_{i};\theta),p(y_{3}|x_{i};\theta),\ldots,p(y_{k}|x_{i};\theta)$ is the probability distribution finally output by the model. We use Kullback-Leibler divergence in model’s training process, the reason and its difference with cross entropy will be discussed thoroughly in Section 4.2. Our goal is to find best parameters $\theta^{*}$ that satisfies:
$\begin{split}\theta^{*}&=\arg\min\limits_{\theta}\sum_{i}KL(D^{{}^{\prime}}_{i}||P_{i})\\ &=\arg\min_{\theta}\sum_{i}\sum_{j}\left(d^{y_{k}}_{x_{i}}\ln\frac{d^{y_{k}}_{x_{i}}}{p(y_{k}|x_{i};\theta)}\right)\end{split}$ (1)
@@
$\theta^{*}=\arg\max_{\theta}\sum_{i}\sum_{j}\left(d^{y_{k}}_{x_{i}}\ln p(y_{k}|x_{i};\theta)\right)$ (2)
In the meantime, we’ll look at the optimization equation in traditional single-label and multi-label methods, they use cross entropy as the main practice:
$\theta^{*}=H(D^{{}^{\prime}}_{i},P_{i})=-\sum_{i}\sum_{k}d^{y_{k}}_{x_{i}}\log p(y_{k}|x_{i};\theta)$ (3)
$\begin{split}\theta^{*}&=\sum_{i}\frac{1}{\lvert Y_{i}\rvert}H(D^{{}^{\prime}}_{i},P_{i})\\ &=-\sum_{i}\sum_{k}\frac{1}{\lvert Y_{i}\rvert}d^{y_{k}}_{x_{i}}\log p(y_{k}|x_{i};\theta)\end{split}$ (4)
Eq.3 and Eq.4 are the optimization functions of single-label and multi-label method respectively. Their corresponding maximum likelihood functions are as Eq.5 and Eq.6:
$\theta^{*}=\arg\max_{\theta}\sum_{i}\ln p(y(x_{i})|x_{i};\theta)$ (5)
$\theta^{*}=\arg\max_{\theta}\sum_{i}\frac{1}{|Y_{i}|}\sum_{y\in Y_{i}}\ln p(y|x_{i};\theta)$ (6)
For single-label problems, where the model only considers one label, it is necessary to compute the entropy of that particular label. In contrast, multi-label tasks require aggregating the entropies of multiple labels. Compare Eq.2, Eq.5 and Eq.6, the simplified maximum likelihood functions differ only from the ranges of labels involved in the calculation(Geng,, 2016). ). Single- label and multi-label learning are special cases of the distribution of LS. LS can be seen as a more general learning technique due to its correlationship with traditional training methods.
3.2 Training Deep learning Models
Now we need to train the deep learning model. The goal of model training is to iteratively optimize the parameter $\theta$ in order to find the optimal parameter $\theta^{*}$ that satisfies:
$\arg\min\limits_{\theta}\sum_{i}KL(D^{{}^{\prime}}_{i}||P_{i})$ (7)
Thus the loss function can be defined as:
$T=-\sum_{i}KL(D^{{}^{\prime}}_{i}||P_{i})$ (8)
In the forward propagation, there is:
$output=forward(input,\theta_{i})$ (9)
Where $input$ is the input data $(x_{i1},x_{i2},x_{i3},\ldots,x_{ik})$, $\theta$ is the parameter value at time step $i$, $output$ is the result of the model. The specific calculation process of forward propagation varies according to the structure of the model.
Then, We use SGD method to optimize the model. In the backward propagation, according to the chain rule , we have:
$\delta_{\text{output}}=\frac{\partial\text{T}}{\partial output}$ (10)
$\delta_{i}=\delta_{i+1}\cdot\frac{\partial a_{i}}{\partial z_{i+1}}\cdot\frac{\partial z_{i}}{\partial a_{i}}$ (11)
For each layer $i$, the gradient is propagated backward from the output layer. $\delta_{i}$ represents the gradient of layer $i$, $\delta_{i+1}$ represents the gradient of the next layer, the input for each layer is $z_{i}$, the output of the activation function for each layer is $a_{i}$.
The SGD algorithm randomly selects a group from the sample, updates it and then extracts another group and updates it again(Adikari and Draper,, 2021). In the case of extremely large sample size, it is possible to obtain a model with a loss value within the acceptable range without training all samples.(Prasher et al.,, 2023; Subramaniam et al.,, 2023) The iteration update formula for SGD (Stochastic Gradient Descent) is:
@@ 3.2 Training Deep learning Models
where $\theta_{i+1}$ is the parameter vector at iteration $i+1$, $\theta_{i}$ is the parameter vector at iteration $i$, $\eta$ is the learning rate, which controls the step size of each update. $\delta_{i}$ is the gradient of the loss function T with respect to the parameter $\theta_{i}$.
3.3 Selected Deep Dearning Architectures
@@
TextCNN, BERT, and RoBERTa are three highly influential and successful deep learning architectures in the field.
@@
TextCNN, as a pioneering architecture, introduced the idea of applying convolutional neural networks to text classification tasks. It leverages convolutional layers with various kernel sizes to capture local features in text data. TextCNN has demonstrated strong performance and efficiency in handling fixed-length input sequences(Wang et al., 2022b, ; Jiang,, 2022).
@@
Further, BERT (Bidirectional Encoder Representations from Transformers) brought about a breakthrough in language understanding. BERT introduced the concept of pre-training a transformer-based model on large amounts of unlabeled text data and fine-tuning it for downstream tasks. It showed remarkable success in a wide range of natural language processing tasks, including question answering, sentiment analysis, and named entity recognition.
@@
RoBERTa, building upon BERT’s foundation, refined the pre-training process and achieved improved performance. It incorporated additional pre-training data and introduced modifications to the training objectives. RoBERTa demonstrated enhanced language representation capabilities and surpassed BERT’s performance on various benchmarks and tasks.
In our experiments, we aim to test the generalizability of LS across different architectures. By evaluating our method on these three models, we aim to showcase the applicability and effectiveness of this approach across diverse deep learning architectures.
(I) TEXTCNN
The structure of TextCNN consists of a convolutional layer, a max pooling layer, and a softmax layer for n-class classification.
@@
Figure 2: Workflow of TextCNN sentiment classification architecture
Firstly, the textual input will undergo convolutional operations using different sizes of convolutional kernels, aiming to capture the textual features(Zhang et al.,, 2020). Assuming the input is a text sequence $x=(x_{i1},x_{i2},x_{i3},\ldots,x_{ik})$, where $x_{i}$ represents the word vector for the $i-th$ word. The convolutional layer applies multiple convolutional filters to extract local features:
$c_{i}=f(W\cdot x_{i:i+h-1}+b)$ (13)
Here, $x_{i:i+h-1}$ represents the concatenation of word vectors from the $i-th$ word to the $(i+h-1)-th$ word, W is the weight matrix of the convolutional filters, b is the bias vector, and $f$ denotes the activation function.
After the convolution is completed, the output will undergo a pooling layer for further processing. The max pooling operation selects the maximum value from each feature map as the pooled feature value:
$p_{i}=\max(c_{i:i+h-1})$ (14)
Here, $c_{i:i+h-1}$ represents the feature vector from the $i-th$ feature to the $(i+h-1)-th$ feature.
Finally, The pooled feature vector serves as the input to the fully connected layer, which performs linear transformation and non-linear activation,:
$y=\text{softmax}(W_{y}\cdot p+b_{y})$ (15)
$W_{y}$ is the weight matrix of the fully connected layer, $b_{y}$ is the bias vector, and $softmax$ is the activation function for multi-class classification.
TextCNN’s simplicity, lower parameter count, and computational efficiency contribute to its fast training speed(Suyan et al.,, 2019). It has demonstrated good performance in text classification tasks.
(II) BERT
BERT (Bidirectional Encoder Representations from Transformers) is a pre-trained language model that utilizes a transformer-based architecture for various natural language processing tasks(Palani et al.,, 2021). The model incorporates both left and right context during training, allowing it to capture deep contextual relationships in text(Sebastian et al.,, 2022).
Figure 3: Workflow of BERT sentiment classification architecture
The input to BERT consists of tokenized text, where each token represents a word. Special tokens like [CLS] and [SEP] are added at the beginning and end of each sentence respectively. Let $H$ be the input sequence $X$ processed by the embedding layer.
Input $H$ is to be split into multiple heads: $H=[h_{1},h_{2},\ldots,h_{h}]$, where each $h$ is a head containing a subset of hidden states.
For each head $h$, we initialize its own set of weight matrix: $Key(K)$, $Value(V)$, and $Query(Q)$. These heads are then linearly transformed and passed through the softmax function to obtain:
$\begin{split}Z_{i}&=\text{Attention}(QW^{Q}_{i},KW^{K}_{i},VW^{V}_{i})\\ &=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V\end{split}$ (16)
The outputs of each head are concatenated to form the multi-head attention output:
$\begin{split}Z&=\text{MultiHead}(Q,K,V)\\ &=\text{Concat}(Z_{1},Z_{2},\ldots,Z_{h})\end{split}$ (17)
This step involves repeating the self-attention operation multiple times, allowing the model to learn information from different perspectives and enrich the semantic representation.
Next, the result $Z=[z_{1},z_{2},\ldots,z_{h}]$ is fed into a feed-forward neural network layer. This layer applies multiple matrix transformations and activation functions introducing non-linearity and obtaining the final vector representation:
$F=relu(W_{2}\cdot(W_{1}Z+b_{1})+b_{2})$ (18)
After the feed-forward neural network layer, a pooling operation is performed to obtain the aggregated representation. The pooling method used in this experiment is max pooling, where the maximum value is selected across each dimension of the representation.
Then, the pooled representation is passed to the classification layer. This layer typically consists of one or more fully connected layers followed by a softmax activation function. The classification layer applies a series of linear transformations and activation functions to the pooled representation, mapping it to corresponding class probabilities.
BERT employs a stack of transformer encoders, composed of self-attention and feed-forward neural network layers. This allows the model to attend to different parts of the input sequence and capture word dependencies. BERT is trained using Masked Language Model(MLM), where a percentage of input tokens are masked, and the model predicts these masked tokens based on context. It also uses Next Sentence Prediction (NSP), where it learns to predict if a second sentence follows the first(Trang and Shcherbakov,, 2020). BERT can be fine-tuned on specific tasks by adding task-specific layers and training on labeled data(Riaz et al.,, 2022). BERT has achieved state-of-the-art results in various NLP tasks, demonstrating its effectiveness in understanding context and semantic relationships in text(Zou et al.,, 2022).
(III) RoBERTa
The RoBERTa model is an improved version built upon the BERT pre-trained model. RoBERTa introduces three key enhancements to the BERT model(Kamath et al.,, 2022). Firstly, it adopts a dynamic masking mechanism, which generates a new masking pattern for each input sequence presented to the model. This allows for greater variation and exploration during training. Secondly, RoBERTa removes the Next Sentence Prediction task, which was present in BERT. By eliminating NSP, the model’s efficiency is improved as it no longer needs to predict the relationship between two sentences(Xu and Hou,, 2022; Choure et al.,, 2022). Lastly, RoBERTa increases the size of the training corpus and adjusts the batch size and training steps accordingly(Mu et al.,, 2021; Qi et al.,, 2022). This expansion of training data and adjustments in batch size and steps contribute to enhanced learning and better model performance.
4 Experimental Results

### 23 Feb 2024 v2

# Revisiting the Role of Label Smoothing in Enhanced Text Sentiment Classification

@@ Abstract
Label smoothing is a widely used technique in various domains, such as text classification, image classification and speech recognition, known for effectively combating model overfitting. However, there is little fine-grained analysis on how label smoothing enhances text sentiment classification. To fill in the gap, this article performs a set of in-depth analyses on eight datasets for text sentiment classification and three deep learning architectures: TextCNN, BERT, and RoBERTa, under two learning schemes: training from scratch and fine-tuning. By tuning the smoothing parameters, we can achieve improved performance on almost all datasets for each model architecture. We further investigate the benefits of label smoothing, finding that label smoothing can accelerate the convergence of deep models and make samples of different labels easily distinguishable.
Index Terms:
Label smoothing, Deep learning, BERT, RoBERTa, Sentiment classification
I Introduction
@@
Text sentiment classification is to identify and extract emotional tendencies (such as positive, negative or neutral) from text, which provides important application scenarios and challenges for the development of natural language processing (NLP) (Li et al., 2022). Many methods have been proposed for this task, for instance, convolutional neural networks (CNN) (Zhang et al., 2015), and recurrent neural networks (RNN) with attention mechanism (Zhang et al., 2016; Chen et al., 2017). Transformer-based models (Jiang et al., 2021; Chang et al., 2020) have excelled in sentiment classification tasks, achieving state-of-the-art performance (Liu et al., 2022a, b).
@@
As a commonly used regularization method to overcome overfitting, label smoothing (LS) proceeds by using soft targets that are a weighted average of the hard targets and the uniform distribution over labels (Müller et al., 2019; Lienen and Hüllermeier, 2021). This technique lessens the disparity between the top probability estimate and the remaining ones, thereby acting as a barrier to the model from generating extremely confident predictions, consequently decreasing the model’s likelihood of becoming excessively tailored to the training data (Lukasik et al., 2020; Gao et al., 2020). This also fosters generalization, leading to better performance on unseen data (Chen et al., 2020; Cui et al., 2022). Furthermore, LS can also mitigate the impact of noisy labels on the training process (Li et al., 2020).
@@
LS has achieved widespread success in NLP (Chandrasegaran et al., 2022; Liu et al., 2022c) by introducing soft targets, which allows the model to optimize towards a more flexible direction during training (Haque et al., 2023; Pan et al., 2022). However, the fine-grained research on LS for text sentiment classification is still limited. For the task of sentiment classification, emotion is not an absolute and discrete concept (Pang et al., 2002; Onan, 2022), but there is certain fuzziness and continuity (Hasib et al., 2021). LS can take account of this ambiguity, allowing the model to learn more balanced relative probabilities between different emotions.
@@
In order to investigate how LS benefits sentiment classification, in this paper we conduct extensive experiments on three widely used deep neural network architectures: TextCNN, BERT, and RoBERTa, under two learning schemes: training from scratch and fine-tuning. Our contributions can be summarized as follows:
  • Through varying the tuning smoothing parameter, LS methods outperform the three baseline architectures on all eight datasets, including six three-class datasets and two binary-class datasets.
  • From in-depth analysis, LS can accelerate the training process of deep models with the deployment of soft labels
  • LS can produce better hidden representations for training examples as they are easier to distinguishable than those produced by the baseline method.
@@
The remaining sections of this paper are organized as follows. We first review the related works in Section 2. Then, in Section 3, we propose the application and deployment of LS in text sentiment classification. Subsequently, in Section 4, we present the experimental results on eight sentiment analysis datasets. Finally, Section 5 provides a discussion and conclusion of this paper.
II Related Works
This paper is related to two lines of research: text sentiment classification and label smoothing.
II-A Text Sentiment Classification
@@
Text sentiment classification is one special type of text classification Kowsari et al. (2019); Si et al. (2020), as the labels are ordinal. In early days, statistical-based models like Naive Bayes (Kim et al., 2006; Raschka, 2014) and support vector machines (Tong and Koller, 2001; Lilleberg et al., 2015) were dominant in text sentiment classification. These models, while accurate and stable, required time-consuming feature design and often overlooked the context information in text data (Kowsari et al., 2019). Since the 2010s, deep learning models (Zulqarnain et al., 2020; Liu et al., 2016), which automatically extract meaningful representations for text without the need for manual rule and feature design (Minaee et al., 2021), are increasingly utilized for sentiment classification.
Although deep learning methods have achieved good performance on text sentiment classification, they may still suffer issues like slow convergence, poor performance. In this paper, we explore how LS benefits deep sentiment methods.
II-B Label Smoothing (LS)
@@
LS is firstly introduced in the field of computer vision (CV) and has achieved success in various visual recognition tasks (Gao et al., 2017; Müller et al., 2019; Si et al., 2022). Later, the method is shown to be effective in Machine Translation (MT) (Gao et al., 2020). Moreover, it also has applications in sentiment classification (Luo et al., 2022, 2021) and Named Entity Recognition(NER) (Zhu and Li, 2022; Yu et al., 2022), improving model calibration and bringing flatter neural minima.
@@
Meanwhile, researchers have started to apply LS to text sentiment classification. One attempt is to add LS to the loss function to enhance the performance of the model (Wang et al., 2022a). Similarly, another group add LS to loss function and perform emotion classification on the adaptive fusion features obtained (Yan et al., 2023). Also, properties of LS and its adversarial variants are studied, showing LS can enhance the adversarial robustness of the model (Yang et al., 2022). Yan et al. (2023) proposes a new cyclic smoothing labeling technique for handling the periodicity of angles and increasing error tolerance for adjacent angles. They also designed a densely-coded tag that greatly reduced the length of the code. Wu et al. (2022) propose an efficient data augmentation method, termed text smoothing, by converting a sentence from its one-hot representation to a controllable smoothed representation and show that text smoothing outperforms various mainstream data augmentation methods by a substantial margin.
@@
Although LS is a common label processing technique, there are still some problems and challenges that need further study. Examples include comparison of effects across different data sets, tasks, and model structures, comparison with other regularization methods, etc (Huang et al., 2020). Studying LS can explore these problems and promote the further development of the field of LS. Our work revisits the application of LS to sentiment classification, and conducts an in-depth analysis of its power.
III Label Smoothing Method for Text Classification
@@
Given a tokenized input $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{n})$ and a set of labels $\boldsymbol{y}=(y_{1},y_{2},\ldots,y_{k})$, where $n$ is the length of input and $k$ is the number of categories for classification. Thus, $\boldsymbol{D}_{i}=(d_{x_{i}}^{y_{1}},d_{x_{i}}^{y_{2}},\ldots,d_{x_{i}}^{y_{k}})$ describes the extent to which the instance $x_{i}$ belongs to a label $y_{k}$. We expect to obtain a final classification result $d_{i}$ which is typically the maximum value of the label distribution $\boldsymbol{D}_{i}$ obtained by applying a softmax normalization to the last layer of the deep neural networks. The label corresponding to the maximum value represents the category to which this document belongs.
III-A Basics of Label Smoothing
@@
Compared with the commonly used hard target, label distribution has some advantages. Due to the continuity of label distribution, it has a larger labeling space, a broader range of expression, and therefore can provide greater flexibility in the learning process. Specifically, LS is a straightforward way to convert hard labels to soft label distributions, which is a mixture of the one-hot hard label and the uniform distribution, i.e., $\boldsymbol{D}_{i}^{{}^{\prime}}=(1-k\lambda)\boldsymbol{D}_{i}+\lambda\boldsymbol{1},\lambda\in[0,1]$, where the original one-hot distribution $\boldsymbol{D}_{i}=(d_{x_{i}}^{y_{1}},d_{x_{i}}^{y_{2}},\ldots,d_{x_{i}}^{y_{k}})$, $\boldsymbol{1}=(1,1,\ldots,1)_{1\times{k}}$, and $\lambda$ is the smoothing parameter.
@@
Suppose that $\theta$ is the trainable parameters of a classification model, and for the $i$-th training example, $P(\boldsymbol{y}|x_{i};\theta)=p(y_{1}|x_{i};\theta),p(y_{2}|x_{i};\theta),\ldots,p(y_{k}|x_{i};\theta)$ is the probability distribution finally output by the model. We use Kullback-Leibler divergence during the training process,that is, the optimal parameter $\theta^{*}$ satisfies:
$\begin{split}\theta^{*}&=\arg\min\limits_{\theta}\sum_{i}KL(\boldsymbol{D}^{{}^{\prime}}_{i}||P_{i})\\ &=\arg\min_{\theta}\sum_{i}\sum_{j}\left(d^{{}^{\prime}y_{j}}_{x_{i}}\ln\frac{d^{{}^{\prime}y_{j}}_{x_{i}}}{p(y_{j}|x_{i};\theta)}\right)\end{split}$ (1)
@@
$\theta^{*}=\arg\max_{\theta}\sum_{i}\sum_{j}\left(d^{{}^{\prime}y_{j}}_{x_{i}}\ln p(y_{j}|x_{i};\theta)\right).$ (2)
For the original hard target multi-classification problem, the optimization equation from the cross-entropy loss is presented:
$\theta^{*}=\arg\min_{\theta}\sum_{i}\sum_{j}d^{y_{j}}_{x_{i}}\log p(d^{y_{j}}_{x_{i}}|x_{i};\theta)$ (3)
The corresponding maximum likelihood function is shown as follows:
$\theta^{*}=\arg\max_{\theta}\sum_{i}\ln p(y(x_{i})|x_{i};\theta),$
where $y(x_{i})$ is the underlying true label for the $i$-t training instance.
For single-label problems, where the model only considers one label, it is necessary to compute the entropy of that particular label. Single-label learning is a special cases of the distribution of LS (Geng, 2016). LS can be seen as a more general learning technique due to its relationship with traditional training methods.
Now we illustrate how we train the deep learning models with LS. The goal of model training is to iteratively optimize the parameter $\theta$ in order to find the optimal parameter $\theta^{*}$ that satisfies:
$\arg\min_{\theta}\sum_{i}KL(\boldsymbol{D}^{{}^{\prime}}_{i}||P(\boldsymbol{y}|x_{i};\theta)$
The stochastic gradient descent (SGD) algorithm and its variants like Adam (Kingma and Ba, 2015) are used to update the parameters. For each mini-batch, the updating formula is
@@
where $\theta_{i+1}$ is the parameter vector at iteration $i+1$, $\theta_{i}$ is the parameter vector at iteration $i$, $\eta$ is the learning rate, which controls the step size of each update. $\delta_{i}$ is the gradient of the loss function with respect to the parameter $\theta_{i}$.
III-B Selected Deep Learning Architectures
@@
TextCNN, BERT, and RoBERTa are three highly influential and successful deep learning architectures in natural language processing.
@@
TextCNN (Kim, 2014), as a pioneering architecture, introduced the idea of applying convolutional neural networks to text classification tasks. It leverages convolutional layers with various kernel sizes to capture local features in text data. TextCNN has demonstrated strong performance and efficiency in handling fixed-length input sequences (Wang et al., 2022b; Jiang, 2022).
@@
BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2018) brought about a breakthrough in language understanding. BERT introduced the concept of pre-training a transformer-based model on large amounts of unlabeled text data and fine-tuning it for downstream tasks. It showed remarkable success in a wide range of natural language processing tasks, including question answering, sentiment analysis, and named entity recognition.
@@
RoBERTa (Liu et al., 2019), building upon BERT’s foundation, refined the pre-training process and achieved improved performance. It incorporated additional pre-training data and introduced modifications to the training objectives. RoBERTa demonstrated enhanced language representation capabilities and surpassed BERT’s performance on various benchmarks and tasks.
In our experiments, we investigate the effectiveness of LS across three aforementioned architectures. The workflow of our method is shown in Fig. 1, taking TextCNN as the example.
@@
Figure 1: Workflow of our sentiment classification with label smoothing. Here the deep architecture is TextCNN, but the workflow for BERT and RoBERTa is similar.
IV Experimental Results and Analysis